Bob paid for his breakfast with 36 coins consisting of dimes and quarters. if the bill was 4.50 then how amny of each type of coin did he use.
Answer by Edwin McCravy(20055)

Bob paid for his breakfast with 36 coins 
consisting of dimes and quarters. if the bill 
was 4.50 then how amny of each type of coin 
did he use.

----------------------------------------------

Two ways to solve.  Use whichever one you are
presently studying:

----------------------------------------------  

Method 1. Using one unknown

Let Q = the number of quarters.

Then to find the number of dimes, use the formula

(SECOND NUMBER) = (TOTAL OF BOTH NUMBERS) — (FIRST NUMBER)

so

NUMBER OF DIMES = 36 - Q

The value of all the quarters = .25Q

The value of all the dimes = .10(36 — Q)

Value of quarters + Value of dimes = $4.50 

                .25Q + .10(36 — Q) = 4.50

Solve that for Q and get Q = 6, then use
 
NUMBER OF DIMES = 36 - Q = 36 - 6 = 30

So the answer is 6 quarters and 20 dimes.

---------------------------------------------

Method 2. Using two unknowns:

          D + Q = 36
    .10D + .25Q = 4.50
  
Solve that system and get D = 30, Q = 6

Edwin